The volume of the cylinder is 324 cm3 and the diameter is 6cm,calculate the height
1
Expert's answer
2019-11-13T11:54:02-0500

The formula for the volume of the cylinder is:

"V =\pi×r^2×h",

hence

"h=\frac{324}{\pi (\frac{6}{2})^2}=11.46" (cm).